OpenStudy (anonymous):

Find the volume of a rectangular solid with the following dimensions: 2√3 cm x 7√5 cm x 4√6 cm.

OpenStudy (anonymous):

multiply \[2\sqrt{3}\times 7\sqrt{5}\times 4\sqrt{6}\]